OpenStudy (anonymous):

if josh can wash a car in 2 hours and edward can do it in 3, how fast can they do it together?

OpenStudy (anonymous):

I think it so: u must consider the hours in 1 minute this mean only this 1 divide by hours in minutes ok u must follow only the steps:

OpenStudy (anonymous):

\[(\frac{ 1 }{ 120 }+\frac{ 1 }{ 180 })*x=1\]

OpenStudy (anonymous):

now I multiply the numbers in parentetes the denominator must be the same: so:

OpenStudy (anonymous):

\[\frac{ 1 }{ 120 }*\frac{ 3 }{ 3 } ~~ then~~ \frac{ 1 }{ 180}*\frac{ 2 }{ 2 } \]

OpenStudy (anonymous):

\[(\frac{ 3 }{ 360}+\frac{ 2 }{ 360})=\frac{ 5 }{ 360 }\]

OpenStudy (anonymous):

Now we have:

OpenStudy (anonymous):

\[(\frac{ 5 }{ 360 })*x=1 \]

OpenStudy (anonymous):

Now I divede both side with 5/360

OpenStudy (anonymous):

it become:

OpenStudy (anonymous):

\[x=1:(\frac{ 5 }{ 360 })\]

OpenStudy (anonymous):

now we have this rule 1/2:1/2 in this case the dived became * and 1/2 became inverted

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*\frac{ 2 }{ 1}=\frac{ 2 }{ 2 }=1\]

OpenStudy (anonymous):

\[x=1*\frac{ 360}{ 5 }=\frac{ 360 }{ 5 }=72\]

OpenStudy (anonymous):

72 minutes.